OpenStudy (anonymous):

could any one help me put this in a table y = –4x + 3?

OpenStudy (anonymous):

u can see that m/gradient=-4 so it would be slopping down for y-intercept u make x=0 y=-4(0) + 3 y=3 (0,3) for x-intercept u make y=0 0=-4x + 3 x=3/4 (3/4,0) plot those points and join it together.

OpenStudy (anonymous):